OpenStudy (anonymous):

what about this one: Which is the largest of the following units of length in the U.S. customary system of measurement?

OpenStudy (anonymous):

answer choices: feet, inches, miles, or yards

OpenStudy (anonymous):

dont show ur work

OpenStudy (anonymous):

ok, this one is pretty easy. The ruler that you use in school, measures in inches. 12 inches is 1 foot. 3 feet is 1 yard. 5,280 feet is 1 mile. Which one do you think it is?

OpenStudy (anonymous):

idk thats why im asking u guys.... but if it was my choice i would say i guess miles

OpenStudy (anonymous):

yes, miles is correct.

OpenStudy (anonymous):

The average hair growth rate is one-half inch per month. Express this rate in millimeters per month. answer choices: 1.27mm/month, 3.81mm/month, 12.7mm/month, or 38.1mm/month

OpenStudy (anonymous):

25.4 mm=1inch so 25.4*.5=12.7mm/month

OpenStudy (anonymous):

Rachel went hiking for one hundred fifty minutes. Express this time in hour and minutes. answer choices: 1hr 5 mins, 1 hr 50 mins, 2hr 5mins, or 2 hr 30 mins

OpenStudy (anonymous):

150 minutes divided by 60mins/hr = 2.5

OpenStudy (anonymous):

so 2hr 5 mins

OpenStudy (anonymous):

sorry, I meant 2.5hrs or 2hrs. 30mins.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A certain type of truck in the trucking industry is legally allowed to have a maximum weight load of 7 tons. Express this weight load in pounds. answer choices: 1,400lb, 3,500lb, 7,000lb, 14,000lb

OpenStudy (anonymous):

dont show work

OpenStudy (anonymous):

2,000 pounds = 1 ton 7*2,000 = 14,000

OpenStudy (anonymous):

On a recipe, it's recommended that the fish filets be cooked until reaching an internal temperature of one hundred forty degrees Fahrenheit. What is the equivalent temperature in degrees Celsius? answer choices: 46degrees Celsius, 60 degrees Celsius, 108 degrees Celsius or 284 degrees Celsius

OpenStudy (anonymous):

i think a but not sure

OpenStudy (anonymous):

140 degrees Fahrenheit - 32 = 108 108 degrees Fahrenheit * 5 = 540 540 degrees Fahrenheit / 9 = 60 degrees Celsius 60 degrees Celsius is your answer

OpenStudy (anonymous):

One serving of water is considered 8 fluid ounces. It is recommended that adults drink 8 servings of water per day. How many quarts of water is this? answer choices: 1qt, 2qt, 4qt, or 8qt

OpenStudy (anonymous):

i think D

OpenStudy (anonymous):

There are 32 ounces in a quart. 8*8 = 64 64/32 = 2 2 quarts

OpenStudy (anonymous):

Brb